# Quantitative Contraction Rates for McKean-Vlasov Stochastic Differential Equations with Multiplicative Noise

### v1

@@ 1 Introduction
We adopt the coupling method in Eberle et al., (2019), which deals with classical SDEs and McKean-Vlasov SDEs with additive noise, to derive quantitative Harris-type theorems for McKean-Vlasov processes with multiplicative noise and general drift coefficients. Using a similar coupling method as that of Eberle et al., (2019) we obtain explicit quantitative contraction rates for the McKean-Vlasov SDE given in (1.1). The essence of our approach uses a mixed coupling. Such a method uses a synchronous coupling when the two marginal processes are close to each other and a reflection coupling when they are farther apart. Such an idea was also used in Zimmer, (2017) for infinite-dimensional SDEs with additive noise. Our first result, Theorem 3.4, shows a contraction in the Wasserstein-1 distance under contractivity at infinity along with certain perturbation bounds on the diffusion coefficient $\sigma$. This result is then used to derive a uniformly in time propagation of chaos bound in Proposition 4.2 and further obtain the existence of a unique invariant measure and exponential ergodicity with an additional growth condition on the drift $b$ in Theorem 4.4. We next present a local contraction result using a multiplicative Wasserstein distance that depends on the initial law under a dissipative drift condition in Theorem 3.7. This contraction result also leads to a uniform in time propagation of chaos bound along with the exponential ergodicity for the process (1.1); see Proposition 4.3 and Theorem 4.5. In Theorems 3.4 and 3.7, the contraction rates can be computed from the coefficients $b$ and $\sigma$ explicitly.
While both Theorems 3.4 and 3.7 are derived using the coupling method, it is worth pointing out the difference between the contraction at infinity and the dissipative drift conditions. Roughly, contraction at infinity pertains to the interplay among the net drift of the marginal processes, which tends to push them closer to each other when they diverge far apart. The dissipative condition, however, ensures the marginal process is steered back towards the origin if it strays too far from it.
@@
[…]
@@ 1.1 Assumptions
$\|\sigma(x)\|\leq M\quad\text{ and }\quad\|\sigma(x)^{-1}\|\leq\Lambda,\qquad\forall x\in\mathbb{R}^{d},$ (1.3)
@@ 2 Preliminary Results
$\mathrm{d}X_{t}$ $=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ $X_{0}=x_{0},$
@@
Let $(x_{0},y_{0})$ and $(B_{t})$ be as before. The reflection coupling of two solutions to (2.1) is a diffusion process $(X_{t},Y_{t})$ with values in $\mathbb{R}^{2d}$ solving
@@
Now we are in a position to construct a coupling for (1.1), which is adapted from that in Eberle et al., (2019). Roughly, the idea is to fix some $\delta\in(0,1)$ and use the reflection coupling when $|X_{t}-Y_{t}|\geq\delta$, and the synchronous coupling when $|X_{t}-Y_{t}|\leq\delta/2$; in between we use a mixture of both couplings.
@@
$\ u_{t}=\begin{cases}\frac{(\sigma(Y_{t}))^{-1}(X_{t}-Y_{t})}{|(\sigma(Y_{t}))^{-1}(X_{t}-Y_{t})|},&\text{ if }X_{t}\neq Y_{t},\\ \text{an arbitrary unit vector},&\text{ if }X_{t}=Y_{t}.\end{cases}$
@@
Consider the coupling $U_{t}=(X_{t},Y_{t})$ from (2.3). Now define the process $Z_{t}=X_{t}-Y_{t}$ and the radial process $r_{t}=|Z_{t}|$. Then we have
@@
For $x\neq y\in\mathbb{R}^{d}$, denoting $z=x-y$, $e=z/|z|$, $u(x,y)=(\sigma(y)^{-1}z)/|\sigma(y)^{-1}z|$, $H(x,y)=I-2u(x,y)u(x,y)^{\top}$, $\Delta(x,y)=\sigma(x)-\sigma(y)$, and $\alpha(x,y)=\sigma(x)-\sigma(y)H(x,y)$. Then
@@
$A:=\sup_{x\neq y}\bigg{\{}\frac{|\sigma(x)-\sigma(y)|^{2}|x-y|^{2}-|(\sigma(x)-\sigma(y))^{\top}(x-y)|^{2}}{|x-y|^{3}}\bigg{\}}.$ (2.10)
Apparently, $A\geq 0$. We also claim that $A<\infty$. Indeed, since $\|\sigma(x)\|\leq M$ thanks to Assumption 1.4, the expression $\frac{|\sigma(x)-\sigma(y)|^{2}|x-y|^{2}-|(\sigma(x)-\sigma(y))^{\top}(x-y)|^{2}}{|x-y|^{3}}$ in (2.10) is bounded when $|x-y|$ is large. On the other hand, Lemma 2.2 ensures that it is also bounded when $|x-y|$ approaches 00. Therefore the claim that $A<\infty$ follows.
@@ 3.1 Contraction Result Under Contractivity at Infinity
Our first result needs the following assumptions.
Assumption 3.1.
The constants $M$ and $\Lambda$ in Assumption 1.4 satisfy
$D:=2\Lambda^{-1}-\sqrt{2}M>\sqrt{2}.$ (3.1)
@@
where $A$ is the nonnegative constant defined in (2.10). Note that $R_{1},R_{2}<\infty$ by (3.2). Next we define
@@
where $\kappa^{*}$ is defined by $\kappa^{*}(r)=\kappa(r)$ when $r\kappa(r)\geq-A$ and $\kappa^{*}(r)=-A/r$ when $r\kappa(r)<-A$. Clearly we have $\kappa^{*}(r)\geq\kappa(r)$ and $r\kappa^{*}(r)+A\geq 0$ for all $r\geq 0$. Moreover, $\varphi(r)=\varphi(R_{1})$ and hence $\Phi(r)=\Phi(R_{1})+(r-R_{1})\varphi(R_{1})$ for $r\geq R_{1}$. Now, we define the constant $c>0$ to be
$c=\begin{cases}\dfrac{D^{2}-2}{2\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s},&\text{ if }2<D^{2}\leq 4,\\[8.61108pt] \displaystyle\frac{1}{\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s},&\text{ if }D^{2}>4.\end{cases}$
Note that
$c\leq\frac{1}{\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s}.$ (3.3)
@@
Note that $\frac{2}{D^{2}}\leq g(r\wedge R_{2})\leq 1$. Finally, the function $f:[0,\infty]\to[0,\infty)$ is given by
@@
[…]
Consequently, $f$ is a strictly increasing and concave function. Note also that $f$ is linear on $[R_{2},\infty)$.
Finally, it is clear that $f(|x-y|)=f(|y-x|)$ and $f(|x-y|)\leq f(|x-z|)+f(|z-y|)$ for all $x,y,z\in\mathbb{R}^{d}$. Thus,
@@
Suppose Assumptions 1.1, 1.2, 1.3, 1.4, 3.1, and 3.2 hold. Let $\mu_{t}$ and $\nu_{t}$ denote the marginal laws of a strong solution $(X_{t})$ of equation (1.1) with initial conditions $X_{0}\sim\mu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ and $X_{0}\sim\nu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$, respectively. Then there exist constants $C$ and $\gamma$ such that
@@
$|\alpha_{s}^{\top}e_{s}|^{2}\geq D^{2}>2.$ (3.12)
Since $\delta<1\leq R_{1}$, it follows that $r_{s}\geq\delta$ if $I_{\{r_{s}\in\{R_{1},R_{2}\}\}}=1$. Moreover, on the set $\{r_{s}\geq\delta\}$, we have $\mathrm{sc}^{2}(U_{s})=0$ and $\mathrm{rc}^{2}(U_{s})=1$ and therefore
$\mathrm{d}[r]_{s}=I_{\{Z_{s}\neq 0\}}[\mathrm{rc}^{2}(U_{s})|\alpha_{s}^{\top}e_{s}|^{2}+\mathrm{sc}^{2}(U_{s})|\Delta_{s}^{\top}e_{t}|^{2}]\mathrm{d}s=|\alpha_{s}^{\top}e_{s}|^{2}\mathrm{d}s\geq D^{2}\mathrm{d}s.$
@@
[…]
@@
Now Assumption 1.1 along with Lemma 3.3 and the fact that $\mathcal{W}_{1}(\mu_{s},\nu_{s})\leq\mathbb{E}[r_{s}]$ yields the bound
@@
where
$\kappa_{r}:=\sup\{|\kappa(s)|:s\in[0,r]\},\quad\forall r>0.$ (3.18)
Since $\mathrm{sc}^{2}(U_{s})+\mathrm{rc}^{2}(U_{s})=1$ and $f^{\prime}(r_{s})\leq 1$, using Lemma 2.2, we get on the set $\{0<r_{s}<\delta\}$
@@
$\frac{\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2}}{2r_{s}}=\frac{\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}}{2r_{s}}\leq\frac{A}{2}.$ (3.20)
Now plugging (3.17), (3.19), and (3.20) into (3.1), we obtain
@@
where we used (3.12), (3.5), (3.6), and Lemma 3.3 to derive the last inequality.
Note that the first term on the right-hand side of (3.1) is zero in expectation. The second term goes to zero as $\delta\to 0$. Therefore if we can show the last term is nonpositive, we will have $\mathbb{E}[\mathcal{E}_{s}]=O(\delta)$. To this end, we have from (3.6) that
@@
[…]
where we used (3.24), (3.23), (3.3), and Lemma 3.3 to derive the last four inequalities. This together with (3.6) imply that
@@
Since $\delta>0$ can be chosen arbitrarily small, (3.8) follows.
@@ 3.2 Contraction Result Under Dissipative Condition
We now replace Assumptions 3.1 and 3.2 with the following two assumptions.
Assumption 3.5.
The constants $M$ and $\Lambda$ in Assumption 1.4 satisfy
$M<\sqrt{2}\Lambda^{-1}.$ (3.27)
@@
Assume assumptions 1.1, 1.2, 1.3, 1.4, 3.5, and 3.6 with $L_{4}\leq L_{1}\leq\lambda/2$. Let $\mu_{t}$ and $\nu_{t}$ denote the marginal laws of a strong solution $(X_{t})$ to (1.1) with initial laws $\mu_{0},\nu_{0}$, respectively. Then there exist a concave, bounded, nondecreasing continuous function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $f(0)=0$, a Lyapunov function $V(x)$ and constants $c,\varepsilon,K_{1},K_{6}\in(0,\infty)$ such that
@@
Note that the Lipschitz coefficient for the law dependence, $L_{1}$, needs to be small enough to obtain the contraction result above. We will construct a function $f$ similar to that from Section 3.1, but we will first need the following lemma.
@@
Let $V(x)=1+|x|^{2}$. Suppose Assumptions 1.1, 1.2, 1.3, and 3.6 hold, and $L_{4}\leq L_{1}\leq\frac{\lambda}{2}$. Then there exists a constant $L\in(0,\infty)$ such that
@@
By Itô’s formula with Assumptions 1.4 and 3.6 and the fact that $L_{4}\leq L_{1}$, we have
@@
$S_{3}$ =$=\{(x,y):V(x)+V(y)\leq\frac{2L}{\lambda}\},$
$S_{4}$ =$=\{(x,y):V(x)+V(y)\leq\frac{8L}{\lambda}\},$
@@
As before, put $z=x-y$ and $r=|z|$. Then we define the function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
@@
$h(r)$ $=\frac{2}{D^{2}}\int_{0}^{r}(s\kappa^{*}(s)+A)\mathrm{d}s+\frac{8M}{D}r,$
@@
$\eta^{-1}$ $=\int_{0}^{R_{4}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,\ \ \ \xi^{-1}=\int_{0}^{R_{3}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s.$ (3.34)
In the above and in the remainder of this subsection, $A$ is defined as in (2.10), $\kappa^{*}(r):=\kappa(r)\vee A$, and $D:=2\Lambda^{-1}-\sqrt{2}M$ as in (3.1). Note $D>0$ by Assumption 3.5. We further define the semi-metric
@@
$=-f^{\prime}(r)\left(\frac{2}{D^{2}}(r\kappa^{*}(r)+A)+\frac{8M}{D}\right)-\frac{\eta}{4}\Phi(r)-\frac{\xi}{4}\Phi(r)I_{\{r<R_{3}\}}$
$\leq-f^{\prime}(r)\left(\frac{2}{D^{2}}(r\kappa(r)+A)+\frac{8M}{D}\right)-\frac{\eta}{4}f(r)-\frac{\xi}{4}f(r)I_{\{r<R_{3}\}},$ (3.38)
@@
$-\frac{f^{\prime}(r_{t})}{2}\left(\frac{2}{D^{2}}(r_{t}\kappa(r_{t})+A)+\frac{8M}{D}\right)\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
@@
$-\frac{8Mf^{\prime}(r_{t})}{2D}I_{\{r_{t}>0\}}\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
@@
where again $\kappa_{\delta}:=\sup\{|\kappa(r)|:r\in[0,\delta]\}$. Note that Lemma 2.2 is used to derive the last inequality with $K_{0}=dL_{2}$ as in (3.19). Also recall the bounds $|\alpha_{t}^{\top}e_{t}|^{2}\geq D^{2}$ and $f(r)\leq r$ derived in (3.12) and (3.36), respectively. Then we can compute
@@
Now, using the definition of $A$ in (2.10) and the bound (3.12), we get

### v2

@@ 1 Introduction
We adopt the coupling method in Eberle et al., (2019), which deals with classical SDEs and McKean-Vlasov SDEs with additive noise, to derive quantitative Harris-type theorems for McKean-Vlasov processes with multiplicative noise and general drift coefficients. Using a similar coupling method as that of Eberle et al., (2019) we obtain explicit quantitative contraction rates for the McKean-Vlasov SDE given in (1.1). The essence of our approach uses a mixed coupling. Such a method uses a synchronous coupling when the two marginal processes are close to each other and a reflection coupling when they are farther apart. Such an idea was also used in Zimmer, (2017) for infinite-dimensional SDEs with additive noise. Our first result, Theorem 3.3, shows a contraction in the Wasserstein-1 distance under contractivity at infinity along with certain perturbation bounds on the diffusion coefficient $\sigma$. This result is then used to derive a uniformly in time propagation of chaos bound in Proposition 4.2 and further obtain the existence of a unique invariant measure and exponential ergodicity with an additional growth condition on the drift $b$ in Theorem 4.4. We next present a local contraction result using a multiplicative Wasserstein distance that depends on the initial law under a dissipative drift condition in Theorem 3.6. This contraction result also leads to a uniform in time propagation of chaos bound along with the exponential ergodicity for the process (1.1); see Proposition 4.3 and Theorem 4.5. In Theorems 3.3 and 3.6, the contraction rates can be computed from the coefficients $b$ and $\sigma$ explicitly.
While both Theorems 3.3 and 3.6 are derived using the coupling method, it is worth pointing out the difference between the contraction at infinity and the dissipative drift conditions. Roughly, contraction at infinity pertains to the interplay among the net drift of the marginal processes, which tends to push them closer to each other when they diverge far apart. The dissipative condition, however, ensures the marginal process is steered back towards the origin if it strays too far from it.
@@
Throughout this paper, we denote by $\langle\cdot,\cdot\rangle$ and $|\cdot|$ the Euclidean inner product and the corresponding norm, respectively. Furthermore, $\|\cdot\|$ is the matrix norm induced by the Euclidean norm. If $(E,\rho)$ is a Polish space, $\mathcal{P}(E)$ is the collection of probability measures on $E$ and $\mathcal{P}_{1}(E)$ is the subset of $\mathcal{P}(E)$ with finite first absolute moment. Furthermore, we write $\mu(|\cdot|):=\int_{\mathbb{R}^{d}}|x|\mu(dx)$ if $\mu\in\mathcal{P}_{1}(\mathbb{R}^{d})$. We define as in Villani, (2009) the Wasserstein distance of two probability measures $\mu,\nu\in\mathcal{P}(E)$ on a Polish metric space $(E,\rho)$ by
@@ 1.1 Assumptions
$\|\sigma(x)-\sigma(y)\|\leq M\quad\text{ and }\quad\|\sigma(x)^{-1}\|\leq\Lambda,\qquad\forall x,y\in\mathbb{R}^{d},$ (1.3)
@@
Assumption 1.5.
The constants $M$ and $\Lambda$ in Assumption 1.4 satisfy
$D:=2\Lambda^{-1}-M>0.$ (1.4)
@@
Remark 1.7.
The function $\kappa$ in assumption 1.1 may take negative values. Our first main result will require that $\kappa$ is negative for large inputs.
Finally, since $\kappa$ is continuous, we define
$\kappa_{r}:=\sup\{|\kappa(s)|:s\in[0,r]\},\quad\forall r>0.$ (1.5)
@@ 2 Preliminary Results
$\mathrm{d}X_{t}$ $=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},\ \ \$$X_{0}=x_{0},$
@@
The reflection coupling of two solutions to (2.1) is a diffusion process $(X_{t},Y_{t})$ with values in $\mathbb{R}^{2d}$ solving
@@
Heuristically, the reflection coupling constructs a ”mirror hyperplane” between the two processes $X_{t}$ and $Y_{t}$. The idea is that the Brownian motion should eventually help the process hit the mirror, at which point the two processes coincide. Once the processes coincide, we can run them together by switching to the synchronous coupling. Notice that this idea will not work in the case where the drift $b$ has measure dependence, as we would require both $X_{t}=Y_{t}$ and $\mu_{t}=\nu_{t}$, which will not be guaranteed in the above construction. Nevertheless, we can still use this idea to construct a combined coupling.
We now construct a coupling for (1.1), which is adapted from that in Eberle et al., (2019). Roughly, the idea is to fix some $\delta\in(0,1)$ and use the reflection coupling when $|X_{t}-Y_{t}|\geq\delta$, and the synchronous coupling when $|X_{t}-Y_{t}|\leq\delta/2$; in between we use a mixture of both couplings.
@@
[…]
@@
Consider the coupling $U_{t}=(X_{t},Y_{t})$ as written in (2.3). Now define the process $Z_{t}=X_{t}-Y_{t}$ and the radial process $r_{t}=|Z_{t}|$. Then we have
@@
For $x\neq y\in\mathbb{R}^{d}$, denoting z=xyz=x-y, $e=z/|z|$, $u(x,y)=(\sigma(y)^{-1}z)/|\sigma(y)^{-1}z|$, $H(x,y)=I-2u(x,y)u(x,y)^{\top}$, $\Delta(x,y)=\sigma(x)-\sigma(y)$, and $\alpha(x,y)=\sigma(x)-\sigma(y)H(x,y)$. Then
@@
$A:=\sup_{x\neq y}\bigg\{\frac{d\|\sigma(x)-\sigma(y)\|^{2}|x-y|^{2}-|(\sigma(x)-\sigma(y))^{\top}(x-y)|^{2}}{|x-y|^{3}}\bigg\}$ (2.10)
where $d$ is the dimension. Apparently, $A\geq 0$. We also claim that $A<\infty$. Indeed, since $\|\sigma(x)-\sigma(y)\|\leq M$ thanks to Assumption 1.4, the expression $\frac{d|\sigma(x)-\sigma(y)|^{2}|x-y|^{2}-|(\sigma(x)-\sigma(y))^{\top}(x-y)|^{2}}{|x-y|^{3}}$ in (2.10) is bounded when |xy||x-y| is large. On the other hand, Lemma 2.2 ensures that it is also bounded when |xy||x-y| approaches 0. Therefore the claim that $A<\infty$ follows.
@@ 3.1 Contraction Result Under Contractivity at Infinity
Our first result needs the following assumption.
@@
where $A$ is the nonnegative constant defined in (2.10) and $D$ is defined in (1.4). Note that $R_{1},R_{2}<\infty$ by (3.1). Next we define
@@
where $\kappa^{*}$ is defined by $\kappa^{*}(r)=\kappa(r)$ when $r\kappa(r)\geq-A$ and $\kappa^{*}(r)=-A/r$ when $r\kappa(r)<-A$. Clearly we have $\kappa^{*}(r)\geq\kappa(r)$ and $r\kappa^{*}(r)+A\geq 0$ for all $r\geq 0$. Hence $\varphi(r)\leq 1$ for $r\geq 0$. Moreover, $\varphi(r)=\varphi(R_{1})$ and hence $\Phi(r)=\Phi(R_{1})+(r-R_{1})\varphi(R_{1})$ for $r\geq R_{1}$. Now, we define the constant $c>0$ to be
$c:=\dfrac{D^{2}}{4\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s}$ (3.2)
@@
Note that $\frac{1}{2}\leq g(r\wedge R_{2})\leq 1$. Finally, the function $f:[0,\infty)\to[0,\infty)$ is given by
@@
$\begin{aligned} f^{\prime\prime}(r)&=-\frac{2f^{\prime}(r)}{D^{2}}(r\kappa^{*}(r)+A)-\frac{2c}{D^{2}}\Phi(r)I_{\{r<R_{2}\}}\\ &\leq-\frac{2f^{\prime}(r)}{D^{2}}(r\kappa^{*}(r)+A)-\frac{2c}{D^{2}}f(r)I_{\{r<R_{2}\}}\leq 0.\end{aligned}$ (3.5)
Consequently, $f$ is a strictly increasing and concave function. Note also that $f$ is linear on $R_{2},\infty)$.
Finally, it is clear that f(|xy|)=f(|yx|)f(|x-y|)=f(|y-x|) and $f(|x-y|)\leq f(|x-z|)+f(|z-y|)$ for all $x,y,z\in\mathbb{R}^{d}$. Thus,
@@
Suppose Assumptions 1.1, 1.2, 1.3, 1.4, 1.5, and 3.1 hold. Let $\mu_{t}$ and $\nu_{t}$ denote the marginal laws of a strong solution $(X_{t})$ of equation (1.1) with initial conditions $X_{0}\sim\mu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ and $X_{0}\sim\nu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$, respectively. Then there exist constants $C$ and $\gamma$ such that
@@
Remark 3.4.
In application, one would hope for $\gamma>0$. This requires $L_{1},$ to be small when $R_{1}$ and $R_{2}$ are large. Such requirements are, in fact, natural. We notice that $\kappa(r)<0$ is needed to ensure recurrence of the radial process inside some ball of radius $R_{2}$. However, if $L_{1}$ is large, one could have multiple distinct stationary solutions to (1.1) inside that ball.
@@
$|\alpha_{s}^{\top}e_{s}|^{2}\geq D^{2}>0.$ (3.11)
Since $\delta<1\leq R_{1}$, it follows that $r_{s}\geq\delta$ if $I_{\{r_{s}\in\{R_{1},R_{2}\}\}}=1$. Moreover, on the set $\{r_{s}\geq\delta\}$, $Z_{s}\neq 0$, and we have $\mathrm{sc}^{2}(U_{s})=0$ and $\mathrm{rc}^{2}(U_{s})=1$ therefore,
$\mathrm{d}[r]_{s}=I_{\{Z_{s}\neq 0\}}[\mathrm{rc}^{2}(U_{s})|\alpha_{s}^{\top}e_{s}|^{2}+\mathrm{sc}^{2}(U_{s})|\Delta_{s}^{\top}e_{s}|^{2}]\mathrm{d}s=|\alpha_{s}^{\top}e_{s}|^{2}\mathrm{d}s\geq D^{2}\mathrm{d}s.$
@@
$=\int_{0}^{t}f^{\prime\prime}(r_{s})\left(\mathrm{rc}^{2}(U_{s})|\alpha_{s}^{\top}e_{s}|^{2}+\mathrm{sc}^{2}(U_{s})|\Delta_{s}^{\top}e_{s}|^{2}\right)\mathrm{d}s\ \text{ a.s.}$
@@
We now attempt to bound $\mathcal{E}_{s}$ so that $\mathbb{E}[\mathcal{E}_{s}]\leq(\delta)$. To that end we will first find bounds for each term and substitute in the choice of $\gamma$. First, assumption 1.1 along with Lemma 3.2 and the fact that $\mathcal{W}_{1}(\mu_{s},\nu_{s})\leq\mathbb{E}[r_{s}]$ yields the bound
@@
where $\kappa_{r}$ is define in (1.5). Since $\mathrm{sc}^{2}(U_{s})+\mathrm{rc}^{2}(U_{s})=1$ and $f^{\prime}(r_{s})\leq 1$, using Lemma 2.2, we get on the set $\{0<r_{s}<\delta\}$
@@
$\frac{\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2}}{2r_{s}}=\frac{\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}}{2r_{s}}\leq\frac{d\|\Delta_{s}\||r_{s}|^{2}-|\Delta^{\top}r_{s}|^{2}}{2r_{s}^{3}}\leq\frac{A}{2}.$ (3.18)
In the above we used the bound $\mathrm{tr}(\Delta\Delta)^{\top}\leq rank(\Delta)|\Delta|^{2}\leq d|\Delta|^{2}$. Now plugging (3.16), (3.17), and (3.18) into (3.1), we obtain
@@
where we used (3.11), (3.18), to derive the last inequality.
Note that the first term on the right-hand side of (3.1) is zero in expectation. The second term goes to zero as $\delta\to 0$. Therefore if we can show the last term is nonpositive, we will have $\mathbb{E}[\mathcal{E}_{s}]\leq O(\delta)$. To this end, we have from (3.5) that
@@
$\leq-\frac{\Phi(r_{s})}{\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}ds}\leq-\frac{4c}{D^{2}}\Phi(r_{s})\leq-\frac{4c}{D^{2}}f(r_{s}),$
where we used (3.22), followed by (3.21), followed by the definition of $c$ (3.2), and finally lemma 3.2 to derive the last four inequalities. This together with (3.5) imply that
@@
Recall that $\gamma$ is independent of the coupling, so it does not depend on $\delta$. Therefore, since $\delta>0$ can be chosen arbitrarily small, (3.7) follows.
@@ 3.2 Contraction Result Under Dissipative Condition
We now replace Assumption 3.1 with the following assumption.
@@
Assume assumptions 1.1, 1.2, 1.3, 1.4, 1.5, and 3.5 with $L_{4}\leq L_{1}\leq\lambda/2$. Let $\mu_{t}$ and $\nu_{t}$ denote the marginal laws of a strong solution $(X_{t})$ to (1.1) with initial laws $\mu_{0},\nu_{0}$, respectively. Then there exist a concave, bounded, nondecreasing continuous function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $f(0)=0$, a Lyapunov function $V(x)$ and constants $c,\varepsilon,K_{1},K_{6}\in(0,\infty)$ such that
@@
Remark 3.7.
Note again the Lipschitz coefficient for the law dependence, $L_{1}$, needs to be small enough. The choice of $c$ utilizes the rate, $\lambda$, at which the process returns to the ball, as well as the rate of convergence inside the ball due to the coupling.
We will construct a function $f$ similar to that from Section 3.1, but we will first need the following lemma.
@@
Let $V(x)=1+|x|^{2}$. Suppose Assumptions 1.1, 1.2, 1.3, and 3.5 hold, and $L_{4}\leq L_{1}\leq\frac{\lambda}{2}$. Then there exists a constant $L\in(0,\infty)$ such that
@@
By Itô’s formula with Assumptions 1.4 and 3.5 and the fact that $L_{4}\leq L_{1}$, we have
@@
$S_{3}$ $=\{(x,y):V(x)+V(y)\leq\frac{2L}{\lambda}\},$
$S_{4}$ $=\{(x,y):V(x)+V(y)\leq\frac{8L}{\lambda}\},$
@@
As before, put z=xyz=x-y and $r=|z|$. Then we define the function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
@@
$h(r)$ $=\frac{2}{D^{2}}\int_{0}^{r}(s\kappa^{*}(s)+A)\mathrm{d}s+\frac{8\tilde{M}}{D}r,$
@@
$\eta^{-1}$ $=\int_{0}^{R_{4}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,\ \ \ \xi^{-1}=\int_{0}^{R_{3}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,\quad\tilde{M}=2M+\|\sigma(0)\|.$ (3.31)
Note that $\|\sigma(x)+\sigma(y)\|\leq\tilde{M}$ for all $x,y$. In the above and in the remainder of this subsection, $A$ is defined as in (2.10), $\kappa^{*}(r):=\kappa(r)\vee A$, and $D:=2\Lambda^{-1}-\sqrt{2}M$ as in (1.4). Note $D>0$ by Assumption 1.5. We further define the semi-metric
@@
$=-f^{\prime}(r)\left(\frac{2}{D^{2}}(r\kappa^{*}(r)+A)+\frac{8\tilde{M}}{D}\right)-\frac{\eta}{4}\Phi(r)-\frac{\xi}{4}\Phi(r)I_{\{r<R_{3}\}}$
$\leq-f^{\prime}(r)\left(\frac{2}{D^{2}}(r\kappa(r)+A)+\frac{8\tilde{M}}{D}\right)-\frac{\eta}{4}f(r)-\frac{\xi}{4}f(r)I_{\{r<R_{3}\}},$ (3.35)
@@
$-\frac{f^{\prime}(r_{t})}{2}\left(\frac{2}{D^{2}}(r_{t}\kappa(r_{t})+A)+\frac{8\tilde{M}}{D}\right)\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
@@
$-\frac{8\tilde{M}f^{\prime}(r_{t})}{2D}I_{\{r_{t}>0\}}\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
@@
where again $\kappa_{\delta}:=\sup\{|\kappa(r)|:r\in[0,\delta]\}$. Note that Lemma 2.2 is used to derive the last inequality with $K_{0}=dL_{2}$ as in (3.17). Also recall the bounds $|\alpha_{t}^{\top}e_{t}|^{2}\geq D^{2}$ and $f(r)\leq r$ derived in (3.11) and (3.33), respectively. Then we can estimate the second and fourth terms of (3.2) by
@@
Now, using the definition of $A$ in (2.10) and the bound (3.11), we notice the third term in (3.2) is